# Tree-graph based construction of quantum spin models with exact ground state

### v1

@@ I Introduction
Exact results are particularly valuable in the study of quantum many-body physics. They can serve to verify analytical predictions and also provide a starting point for approximation theories. The most famous example of exact results for the quantum many-body problem is the Majundar-Ghosh modelMajumdar and Ghosh (1969); Majumdar (1970) for the spin-1/2 system on a zigzag ladder lattice. It has been shown that the Majundar-Ghosh model is expressed as a sum of projection operators and has a dimer product state as its doubly degenerate ground states. The Affleck-Kennedy-Lieb-Tasaki modelAffleck et al. (1987) is another famous example for which the exact ground state is obtained. Other models where the exact ground-state solution has been established include the Shastry-Sutherland modelSriram Shastry and Sutherland (1981); Verkholyak et al. (2014), the Richardson modelDukelsky et al. (2004); von Delft and Ralph (2001), and various quantum spin systems.Nakano and Takahashi (1995, 1997); Takano et al. (1996); Verkholyak and Strečka (2013); Morita and Shibata (2016); Bose and Ghosh (1997); Schmidt and Laad (2010); Xian (1995); Verkholyak and Strečka (2012); Hikihara et al. (2017)
@@
where ${\bm{s}}_{i}$ is the $S=1/2$ spin operator at the $i$th site. We suppose that the number of spins in the model, $N$, is even. The exchange interaction constants $J_{ij}$ are assumed to be non-negative, i.e., $J_{ij}\geq 0$. The spatial configuration of $\{J_{ij}\}$ which has a one-to-one correspondence with a given tree graph is determined by the protocol we propose. We then show that when the exchange interaction constants $\{J_{ij}\}$ satisfy a specific condition, the model (1) possesses a singlet-dimer product state as its unique ground state. We also demonstrate that by relaxing the condition on $\{J_{ij}\}$, one can construct a model with the macroscopically degenerate ground states.
@@ II Model and Tree graph
Figure 1: (a) Example of a generic binary tree graph for an eight-spin system. Squares and circles denote vertices and bare $S=1/2$ spins, respectively. The grey area represents the spin block assigned to the vertex $v$ indicated in the panel. (b) A vertex $v$ and its child vertices $c_{\rm L}(v)$ and $c_{\rm R}(v)$. The block spins $\tilde{S}_{c_{\rm L}(v)}$ and $\tilde{S}_{c_{\rm R}(v)}$ are coupled via an exchange interaction with the exchange constant $J_{v}$.
@@
Figure 2: Correspondence between tree graphs and quantum spin models. (a) Perfect-Binary Tree (PBT) with two layers, (b) PBT with three layers, and (c) a generic tree graph. In (a), (b), and (c), squares and circles represent the vertices and bare spins, respectively. (d), (e), and (f) depict the quantum spin models corresponding to (a), (b), and (c), respectively. In (d), (e), and (f), circles and lines denote the $S=1/2$ spins and exchange interactions, respectively.
@@ III.1 Perfect-Binary Tree
$=$ $J_{v2}(\tilde{\bm{S}}_{v_{2}})^{2}+(J_{v_{1{\rm L}}}-J_{v2})(\tilde{\bm{S}}_{v_{1{\rm L}}})^{2}$
$~~~~~~+(J_{v_{1{\rm R}}}-J_{v2})(\tilde{\bm{S}}_{v_{1{\rm R}}})^{2}+{\rm const.}.$
Note that $\tilde{\bm{S}}_{v_{2}}={\bm{s}}_{1}+{\bm{s}}_{2}+{\bm{s}}_{3}+{\bm{s}}_{4}$, $\tilde{\bm{S}}_{v_{1{\rm L}}}={\bm{s}}_{1}+{\bm{s}}_{2}$, and $\tilde{\bm{S}}_{v_{1{\rm R}}}={\bm{s}}_{3}+{\bm{s}}_{4}$. Therefore, if $J_{v_{1{\rm L}}}>J_{v_{2}}$ and $J_{v_{1{\rm R}}}>J_{v_{2}}$, one can immediately find that this block Hamiltonian $\mathcal{H}_{v_{2}}$ has the unique ground state of total spin zero, that is the product state of two singlets on the spin pairs $\{{\bm{s}}_{1},{\bm{s}}_{2}\}$ and $\{{\bm{s}}_{3},{\bm{s}}_{4}\}$. Here, it is also noteworthy that a Hamiltonian $\mathcal{H}_{v_{2}}-J_{p}(\tilde{\bm{S}}_{v_{2}})^{2}$ with a certain $J_{p}$ has the same singlet ground state as that of $\mathcal{H}_{v_{2}}$ as long as $J_{v2}>J_{p}$ holds.
@@ III.2 Generic Tree Graph
Here, $J_{p(v)}$ is the exchange interaction constant allocated to the parent vertex of $v$, and $J_{p(v)}=0$ for the root vertex. From the expression of Eq. (10), it immediately follows that if the condition that the exchange constant at a vertex $v$ is larger than that of its parent vertex, $J_{v}>J_{p(v)}$, holds for every $v$, the magnitude of the block spin $\tilde{S}_{v}$ in the ground state takes the smallest value, that is $\tilde{S}_{v}=0$ ($1/2$) for the block with an even (odd) number of spins.
Figure 4: Correspondence between tree graphs and the ground states. (a) Tree graph containing only vertices whose spin block consists of an even number of spins. (b) Tree graph containing vertices whose spin block consists of an odd number of spins. The numbers indicated for vertices denote the magnitude of the block spin $\tilde{S}_{v}$ for each vertex in the ground state. Red lines represent the spin pairs forming the singlet state in the ground state.
@@
Next, consider the tree graphs containing internal vertices whose spin block is composed of an odd number of spins. Here, let us closely look at the ground state of such a spin block with an odd number of spins. The smallest such block is the three-spin block shown in Fig. 5(a). The ground state of its block Hamiltonian is the product state of the $S=1/2$ doublet of the bare spin ${\bm{s}}_{1}$ and the singlet state formed by the spins ${\bm{s}}_{2}$ and ${\bm{s}}_{3}$. It should be noted that the $S=1/2$ degree of freedom of the spin block stems solely from the bare spin ${\bm{s}}_{1}$. Then, consider the case that two spin blocks are merged at a vertex $v$, where one block contains an odd number of spins and has a single $\tilde{S}_{c_{\rm L}(v)}=1/2$ doublet as the unique ground-state doublet, while the other is the block of an even number of spins with the unique singlet ground state, $\tilde{S}_{c_{\rm R}(v)}=0$ [see Fig. 5(b)]. The ground state of such a merged block is a single $\tilde{S}_{v}=1/2$ doublet, which is the product of the ground states of the individual blocks. Therefore, the ground state of a spin block containing an odd number of spins is a single $\tilde{S}_{v}=1/2$ doublet, which can be traced back to a single bare spin in the block. Such a ground-state doublet survives even when the block is merged with a block with an even number of spins. Then, when the spin block is merged with the other block with an odd number of spins [Fig. 5(c)], the merged spin block has the singlet ground state in which two bare spins representing the ground-state doublet of each block form a singlet pair. We thereby find that the model corresponding to the tree graph considered here has a product state of singlet dimers Eq. (8) as its unique ground state. The pattern of the singlet-dimer distribution depends on the tree graph; Fig. 4 presents examples.
@@ IV Summary
In summary, we have introduced a protocol for constructing an $S=1/2$ antiferromagnetic Heisenberg model with the exact ground state based on a tree graph. The protocol determines the spatial configuration of exchange interactions of the model, which has a one-to-one correspondence with the tree graph. Then, we have proven that when the exchange interactions satisfy the condition that their strength increases as one moves from the root vertex to the boundary of the tree graph, the model possesses a product state of singlet dimers as its unique ground state. It has also been shown that one can construct a model possessing the massively degenerate ground states by relaxing the condition of the strength of the interactions.
@@ Acknowledgements.
We thank Kouichi Okunishi and Tomotoshi Nishino for fruitful discussions. T.H. was supported by JSPS KAKENHI Grant Number JP24K06881.

### v3

@@ I Introduction
Exact results are particularly valuable in the study of quantum many-body physics. They can serve to verify analytical predictions and also provide a starting point for approximation theories. The most famous example of exact results for the quantum many-body problem is the Majundar-Ghosh modelMajumdar and Ghosh (1969); Majumdar (1970) for the spin-1/2 system on a zigzag ladder lattice. It has been shown that the Majundar-Ghosh model is expressed as a sum of projection operators and has dimer product states as its doubly degenerate ground states. Other famous models where the exact ground-state solution has been established include the Affleck-Kennedy-Lieb-Tasaki modelAffleck et al. (1987), the Shastry-Sutherland modelSriram Shastry and Sutherland (1981); Verkholyak et al. (2014), and the Richardson modelDukelsky et al. (2004); von Delft and Ralph (2001). Additionally, various quantum spin systems with unique, finitely-degenerate, or infinitely-degenerate exact ground states have been proposed.Nakano and Takahashi (1995, 1997); Takano et al. (1996); Verkholyak and Strečka (2013); Morita and Shibata (2016); Bose and Ghosh (1997); Schmidt and Laad (2010); Xian (1995); Verkholyak and Strečka (2012); Hikihara et al. (2017); Takano (1994); Kumar (2002)
@@
where ${\bm{s}}_{i}$ is the $S=1/2$ spin operator at the $i$th site. We suppose that the number of spins in the model, $N$, is even. The exchange interaction constants $J_{ij}$ are assumed to be positive, i.e., $J_{ij}>0$. The spatial configuration of $\{J_{ij}\}$ which has a one-to-one correspondence with a given tree graph is determined by the protocol we propose. We then show that when the exchange interaction constants $\{J_{ij}\}$ satisfy a specific condition, the model (1) possesses a singlet-dimer product state as its unique ground state. We also demonstrate that by relaxing the condition on $\{J_{ij}\}$, one can construct a model with the macroscopically degenerate ground states.
@@ II Model and Tree Graph
Figure 1: (a) Example of a generic binary tree graph for an eight-spin system. Squares and circles denote vertices and bare $S=1/2$ spins, respectively. The grey area represents the spin block assigned to the vertex $v$ indicated in the panel. (b) A vertex $v$ and its child vertices $c_{\rm L}(v)$ and $c_{\rm R}(v)$. The block spins $\tilde{S}_{c_{\rm L}(v)}$ and $\tilde{S}_{c_{\rm R}(v)}$ are coupled via an exchange interaction with the exchange constant $2J_{v}$.
@@
Figure 2: (Color online) Correspondence between tree graphs and quantum spin models. (a) Perfect-Binary Tree (PBT) with two layers, (b) PBT with three layers, and (c) a generic tree graph. In (a), (b), and (c), squares and circles represent the vertices and bare spins, respectively. (d), (e), and (f) depict the quantum spin models corresponding to (a), (b), and (c), respectively. In (d), (e), and (f), circles and lines denote the $S=1/2$ spins and exchange interactions, respectively.
@@ III.1 Perfect-Binary Tree
$=$ $J_{v_{2}}(\tilde{\bm{S}}_{v_{2}})^{2}+(J_{v_{1{\rm L}}}-J_{v_{2}})(\tilde{\bm{S}}_{v_{1{\rm L}}})^{2}$
$~~~~~~+(J_{v_{1{\rm R}}}-J_{v_{2}})(\tilde{\bm{S}}_{v_{1{\rm R}}})^{2}+{\rm const.}.$
Note that $\tilde{\bm{S}}_{v_{2}}={\bm{s}}_{1}+{\bm{s}}_{2}+{\bm{s}}_{3}+{\bm{s}}_{4}$, $\tilde{\bm{S}}_{v_{1{\rm L}}}={\bm{s}}_{1}+{\bm{s}}_{2}$, and $\tilde{\bm{S}}_{v_{1{\rm R}}}={\bm{s}}_{3}+{\bm{s}}_{4}$. Therefore, if $J_{v_{1{\rm L}}}>J_{v_{2}}$ and $J_{v_{1{\rm R}}}>J_{v_{2}}$, one can immediately find that this block Hamiltonian $\mathcal{H}_{v_{2}}$ has the unique ground state of total spin zero, that is the product state of two singlets on the spin pairs $\{{\bm{s}}_{1},{\bm{s}}_{2}\}$ and $\{{\bm{s}}_{3},{\bm{s}}_{4}\}$. Here, it is also noteworthy that a Hamiltonian $\mathcal{H}_{v_{2}}-J_{p}(\tilde{\bm{S}}_{v_{2}})^{2}$ with a certain $J_{p}$ has the same singlet ground state as that of $\mathcal{H}_{v_{2}}$ as long as $J_{v_{2}}>J_{p}$ holds.
@@ III.2 Generic Tree Graph
Here, $J_{p(v)}$ is a half of the exchange interaction constant allocated to the parent vertex of $v$, and $J_{p(v)}=0$ for the root vertex. From the expression of Eq. (10), it immediately follows that if the condition that the exchange constant at a vertex $v$ is larger than that of its parent vertex, $J_{v}>J_{p(v)}$, holds for every $v$, the magnitude of the block spin $\tilde{S}_{v}$ in the ground state takes the smallest value, that is $\tilde{S}_{v}=0$ ($1/2$) for the block with an even (odd) number of spins.
Figure 4: (Color online) Correspondence between tree graphs and the ground states. (a) Tree graph containing only vertices whose spin block consists of an even number of spins. (b) Tree graph containing vertices whose spin block consists of an odd number of spins. The numbers indicated for vertices denote the magnitude of the block spin $\tilde{S}_{v}$ for each vertex in the ground state. Red lines represent the spin pairs forming the singlet state in the ground state.
@@
Next, consider the tree graphs containing internal vertices whose spin block is composed of an odd number of spins. Here, let us closely look at the ground state of such a spin block with an odd number of spins. The smallest such block is the three-spin block shown in Fig. 5(a). The ground state of its block Hamiltonian is the product state of the $S=1/2$ doublet of the bare spin ${\bm{s}}_{1}$ and the singlet state formed by the spins ${\bm{s}}_{2}$ and ${\bm{s}}_{3}$. It should be noted that the $S=1/2$ degree of freedom of the spin block stems solely from the bare spin ${\bm{s}}_{1}$. Then, consider the case that two spin blocks are merged at a vertex $v$, where one block contains an odd number of spins and has a single $\tilde{S}_{c_{\rm L}(v)}=1/2$ doublet as the unique ground-state doublet, while the other is the block of an even number of spins having the unique singlet ground state with $\tilde{S}_{c_{\rm R}(v)}=0$ [see Fig. 5(b)]. The ground state of such a merged block is a single $\tilde{S}_{v}=1/2$ doublet, which is the product of the ground states of the individual blocks. Therefore, the ground state of a spin block containing an odd number of spins is a single $\tilde{S}_{v}=1/2$ doublet, which can be traced back to a single bare spin in the block. Such a ground-state doublet survives even when the block is merged with a block with an even number of spins. Then, when the spin block is merged with the other block with an odd number of spins [Fig. 5(c)], the merged spin block has the singlet ground state in which two bare spins representing the ground-state doublet of each block form a singlet pair. We thereby find that the model corresponding to the tree graph considered here has a product state of singlet dimers Eq. (8) as its unique ground state. The pattern of the singlet-dimer distribution depends on the tree graph; Fig. 4 presents examples.
@@ IV Summary
In summary, we have introduced a protocol for constructing an $S=1/2$ antiferromagnetic Heisenberg model with the exact ground state based on a tree graph. The protocol determines the spatial configuration of exchange interactions of the model, which has a one-to-one correspondence with the tree graph. Then, we have proven that when the exchange interactions allocated to vertices satisfy the condition that their strength increases as one moves from the root vertex to the boundary of the tree graph, the model possesses a product state of singlet dimers as its unique ground state. It has also been shown that one can construct a model possessing the massively degenerate ground states by relaxing the condition of the strength of the interactions.
@@ Acknowledgements.
We thank Kouichi Okunishi and Tomotoshi Nishino for fruitful discussions. T.H. was supported by JSPS KAKENHI Grant Number JP24K06881. hikihara@gunma-u.ac.jp